OpenStudy (anonymous):

Write an equation for the following transformation of y=x: a vertical stretch by a factor of 4.

OpenStudy (anonymous):

if you could explain vertical stretch I could probably help you out.

OpenStudy (anonymous):

Multiplies all y-values of a function by the same factor greater than 1

OpenStudy (anonymous):

For a function f(x) and a constant A, y=af(x) is a vertical stretch when A is greater than 1.

OpenStudy (yttrium):

It's really clear in the meaning the definition of a stretch. Therefore, just substitute 4 into a and you'll get the new equation. :D

OpenStudy (anonymous):

So, y=4f(x) ?

OpenStudy (yttrium):

Yes. And what is f(x)?

OpenStudy (anonymous):

I don't know, =\ that's all the question said. I thought y=4f(x) was the answer but it seemed too simple.

OpenStudy (yttrium):

y = f(x) right this means that y was represented as a function of x. Say for instance we have y = x^2+3. Our f(x), then is x^2+3. Now, do you get the pattern?

OpenStudy (anonymous):

Kinda, but I don't have any other numbers. Are you saying f(x) is 4?

OpenStudy (yttrium):

No. 4 is a. and f(x) is x since it'y y = x

OpenStudy (anonymous):

So, y= 4x?

OpenStudy (yttrium):

Yes! :)

OpenStudy (anonymous):

Thank you!

OpenStudy (yttrium):

No problem. :) Hope you understand the concept.